# Quantum computing with atomic qubit arrays: confronting the cost of connectivity

### v1

@@ VI Outlook
In parallel with the progress on qubit count scaling, measurement fidelity, and gate fidelity there has been substantial recent work on HLL challenges. The heating challenge is common to both trapped ion and neutral atom qubits. In the trapped ion platform it is possible to use sympathetic cooling between two species[204] without affecting the quantum state encoded in internal degrees of freedom. A related idea has been proposed for alkaline earth like and alkali atoms based on state insensitive Rydberg interactions[205], as well as direct laser cooling in a regime that preserves the nuclear spin state[206]. Another, and simpler approach is to invoke a two-species architecture as discussed in Sec. V. Provided the two species couple to very different optical wavelengths they can be independently cooled, while preserving quantum coherence in the other species as has been shown in [193]. Two-species maps naturally onto a logical qubit architecture with one-species acting as data and the other as ancillae. Swapping quantum information back and forth between the species, the other element can be periodically recooled which will allow for operation of deep circuits.
Leakage outside of the computational basis occurs primarily due to polarization errors on gate control beams, and decay from electronically excited and Rydberg states during gate operations. There are approaches for repumping leakage into a depolarization error[207, 208]. To the extent that leakage can be detected or converted into loss at a known array location the leakage becomes an erasure error which can be used to significantly improve the performance of error correcting codes[209, 70, 210, 121]. In the limit where a large fraction of all errors are erasures, and the erasures are biased which means they arise from leakage from only one of the computational qubit states, the code threshold can be increased by close to an order of magnitude[211].
Lastly there is the problem of atom loss. One source of atom loss is collisions with untrapped background molecules. This loss mechanism can be reduced using cryogenic pumping to achieve lower pressure and single atom lifetimes reaching 6000 s have been demonstrated[212, 52, 213]. With good vacuum and negligible background loss there will still be loss due to atoms being left in Rydberg states during gate operations. Loss may be partially mitigated by design of codes that can tolerate missing qubits[214, 215, 216, 217]. Nevertheless a scalable system capable of running long calculations will need to incorporate a mechanism into the control system for monitoring and replacing lost atoms. First demonstrations of this capability that are able to maintain 1000 qubit scale arrays continuously occupied have been demonstrated[54, 53]. Fully integrating loss detection and atom replacement into the error correction cycle remains an outstanding challenge.
@@
I would like to thank my research group at UWM and colleagues at Infleqtion for many stimulating discussions, Matt Otten for comments on the manuscript, and Sam Norrell for technical assistance. The work in Madison on neutral atom quantum computing is generously supported by the US National Science Foundation, the Department of Energy, the Laboratory for Physical Sciences, the Army Research Office, the Wisconsin Alumni Research Foundation, and Infleqtion, Inc. .

### v3

@@ 6 Outlook
In parallel with the progress on qubit count scaling, measurement fidelity, and gate fidelity there has been substantial recent work on HLL challenges. The heating challenge is common to both trapped ion and neutral atom qubits. In the trapped ion platform it is possible to use sympathetic cooling between two species Barrett2003 without affecting the quantum state encoded in internal degrees of freedom. A related idea has been proposed for alkaline earth like and alkali atoms based on state insensitive Rydberg interactions Belyansky2019 , as well as direct laser cooling in a regime that preserves the nuclear spin state Reichenbach2007 . Another, and simpler approach is to invoke a two-species architecture as discussed in Sec. 5. Provided the two species couple to very different optical wavelengths they can be independently cooled, while preserving quantum coherence in the other species as has been shown in Singh2023 . Two-species maps naturally onto a logical qubit architecture with one-species acting as data and the other as ancillae. Swapping quantum information back and forth between the species, the other element can be periodically recooled which will allow for operation of deep circuits.
Leakage outside of the computational basis occurs primarily due to polarization errors on gate control beams, and decay from electronically excited and Rydberg states during gate operations. There are approaches for repumping leakage into a depolarization error Carr2013 ; ICong2022 . To the extent that leakage can be detected or converted into loss at a known array location the leakage becomes an erasure error which can be used to significantly improve the performance of error correcting codes YWu2022 ; SMa2023 ; Scholl2023 ; MChow2024 . In the limit where a large fraction of all errors are erasures, and the erasures are biased which means they arise from leakage from only one of the computational qubit states, the code threshold can be increased by close to an order of magnitude Sahay2023 .
Lastly there is the problem of atom loss. One source of atom loss is collisions with untrapped background molecules. This loss mechanism can be reduced using cryogenic pumping to achieve lower pressure and single atom lifetimes reaching 6000 s have been demonstrated Schymik2021 ; Pichard2024 ; ZZhang2025 . With good vacuum and negligible background loss there will still be loss due to atoms being left in Rydberg states during gate operations. Loss may be partially mitigated by design of codes that can tolerate missing qubits Vala2005b ; Baker2021 ; Perrin2025 ; ZWei2025 . Nevertheless a scalable system capable of running long calculations will need to incorporate a mechanism into the control system for monitoring and replacing lost atoms. First demonstrations of this capability that are able to maintain 1000 qubit scale arrays continuously occupied have been demonstrated Gyger2024 ; Norcia2024 ; NCChiu2025 . Fully integrating loss detection and atom replacement into the error correction cycle remains an outstanding challenge.
@@
I would like to thank my research group at UWM and colleagues at Infleqtion for many stimulating discussions, Matt Otten for comments on the manuscript, and Sam Norrell for technical assistance. The work in Madison on neutral atom quantum computing is generously supported by the US National Science Foundation, the Department of Energy, the Laboratory for Physical Sciences, the Army Research Office, the Wisconsin Alumni Research Foundation, and Infleqtion.